# Agent-Based Insight into Eco-Choices: Simulating the Fast Fashion Shift

### v1

@@ A-B1 Polarized agentset - peer interaction function
If the threshold is set to $m$, then opinions within a range of ±mplus-or-$\pm m$ from an agent’s opinion will have a homogenizing effect. However, opinions beyond this threshold will have an opposite effect, pushing the agent’s opinion toward the opposite end of the spectrum. For instance, if an agent’s opinion is $0.5$ and their friend’s opinion is $0.9$, with a tolerance level of $m=0.2$, the interaction with that friend will contribute to the opinion change of $1-0.9=0.1$, pulling the agent toward the opposite opinion of that of their neighbors.
A-B2 Social media and government feedback loops
The function for social media feedback loop in Eq. (3) is depicted in Fig. (9) and it returns a more extreme value based on the input, representing the influence of social media feedback on an individual’s opinion. The x-axis represents the agent’s current opinion, and the y-axis represents the adjusted opinion that social media will show on agent’s feed. If the output exceeds $1$, it is manually capped at $0.95$ to prevent extreme values. Similarly, if the output is negative, it is adjusted to $0.05$ to avoid values outside the defined range. Although the function is uniform for all agents, the output varies for each agent’s opinion, reflecting the individual calibration of social media platforms.
@@ A-B2 Social media and government feedback loops
Figure 9: Social media feedback loop Eq. (3)
@@
Eq. (4) ensures that an agent’s opinion changes gradually and cannot be drastically affected in a single time step. Let us delve deeper into Eq. (4). Typically, a linear equation is used to adjust an agent’s opinion based on an influence. However, we devised a more sophisticated equation to ensure that an agent’s opinion cannot change drastically in just one step.
@@
Consider an agent with a susceptibility to social media of $0.5$. Note that the function above maintains the same constant susceptibility, $S_{sm}$, throughout the simulation. The two inputs that are changing are the agent’s current opinion and the social media “influence” direction, calculated one step prior using the function described in Eq. (3). Therefore, we have two inputs. To represent this in a 2D graph, we set the agent’s opinion as a parameter, marked as the “Intersection” point on the following graphs, and the x-axis represents the output of the social media function, indicating the opinion it promotes. The y-axis represents the agent’s updated opinion. The graphs for susceptibility level 0.5 and for agent $A_{t-1}$ values of 0, 0.5, and 1 are Fig. (10), Fig. (11), and Fig. (12), respectively.
Figure 10: Social media influence function for susceptibility level of 0.5 and agent’s opinion of 0
Figure 11: Social media influence function for susceptibility level of 0.5 and agent’s opinion of 0.5
Figure 12: Social media influence function for susceptibility level of 0.5 and agent’s opinion of 1
Let us examine the first case of Fid. (10), where the intersection is marked at zero. This graph represents the scenario where the agent’s susceptibility is 0.5 and their opinion at time $t-1$ is zero. The blue line represents Equation (4), while the red line depicts the linear function typically used in such cases and is included for reference. Note that with the red function, the agent’s opinion can change from 0 to 0.4 in just one time step, which is evidently too drastic. However, with the blue function, this is not the case, ensuring a smoother and slower transition. The further away from the original opinion - the smaller the derivative value of the blue line.
In tFig. (11), we have the same scenario but with the agent’s opinion set at 0.5. The intersection point lies on the $y=x$ line, indicating that if the agent’s opinion matches that promoted by social media, there will be no change, and the same value will be returned. Similarly, Fig. (12) illustrates the case where the agent’s opinion is set at 1. Here, the blue line depicts the potential return, while the red line serves as a reference. In all three cases, the blue line represents the values that the agent’s opinion will take based on the input of social media influence. Notice that the derivative of the blue function decreases the further away we get from the intersection point, representing the smoother transition of opinion.
Now, let’s observe how this function changes as the susceptibility level varies. In Fig. (13), we plot two cases: on the left, where the agent’s susceptibility is 0, and on the right, where it is 1. When the susceptibility level is zero, the input becomes irrelevant; social media will not influence the agent’s opinion, thus returning the same value as $A_{t-1}$, which acts as a parameter for the plot. Similarly, for a susceptibility of 1, the function becomes a linear equation, as the agent is willing to undergo drastic opinion changes in a single step.
@@
Figure 13: Social media influence function for susceptibility level of 0 and 1.
A-C Simulations
A-C1 Time management
@@ A-C3 Pseudocode
A-D Figures
Figure 14: Changes in average values for Model (A1) Source: our simulations.
Figure 15: Changes in average values for Model (A2) Source: our simulations.
Figure 17: Changes in average values for Model (A3) Source: our simulations.
Figure 20: Variances of average values for Model (B3) Source: our simulations.

### v2

@@ A.2.1 Polarized agentset - peer interaction function
If the threshold is set to $\tau=m$, then opinions within a range of $\pm m$ from an agent’s opinion will have a homogenizing effect. However, opinions beyond this threshold will have an opposite effect, pushing the agent’s opinion toward the opposite end of the spectrum. For instance, if an agent’s opinion is $0.5$, with a tolerance level of $\tau=0.2$, then the agent will “accept” an opinion within the range $[0.3,0.7]$, but anything that falls outside of that opinion will have the opposite effect. In particular, it will make incentivize the agent to adopt beliefs opposite to that of the agent they are interacting with. Since opinions are defined in the range $[0,1]$, for a given opinion $b\in[0,1]$, we define the opposite opinion as 1b1-b such. Thus, if an agent $i$ with opinion $b_{i}=0.3$ interacts with an agent $j$ with opinion $b_{j}=0.9$ while the tolerance threshold is $\tau=0.2$, then the $j$th agent’s opinion will contribute $1-b_{j}=0.1$ to $i$th agent’s opinion on the topic. The role of $\tau$ is similar to the role of Broadening Behavior introduced in ([7]), in a sense that it introduces a kind of repulsion towards the opinion of others, especially those an agent disagrees with, and encourages agents to diverge. However, it is important to note that the tolerance threshold we introduce is different from the broadening behavior since it does not encourage the agent to be different from their neighbors, but rather to differ more from the group the agent already differs from.
A.2.2 Social media and government feedback loops
Our feedback loop equation for social media, presented in Eq. (3), is designed to simulate real-world algorithms that customize content for users based on their past preferences. This function also incorporates each agent’s susceptibility. Fig.(A.1) illustrates three variations of the social media feedback loop function, each reflecting different levels of susceptibility. In Fig. (1(a)), the susceptibility is set at $S_{sm}=0.2$, indicating a low susceptibility, which results in a wider function. This means the content shown to the agent is similar to what they currently view, with minimal deviation. Fig. (1(b)) depicts a susceptibility of $S_{sm}=0.5$, leading to a steeper function and a greater divergence from the agent’s existing content. Finally, Fig. (1(c)) shows a susceptibility of $S_{sm}=0.8$, resulting in the steepest function of the three, meaning the content presented is significantly more extreme compared to what the agent’s current view. In any case, if the output exceeds $1$, it is manually capped at $0.95$ to prevent extreme values. Similarly, if the output is negative, it is adjusted to $0.05$ to avoid values outside the defined range.
Moreover, the function differs based on the value of $\beta$, and this adjustment is made for all agents. In Fig. (A.2), we see the same feedback loop function for $S_{sm}=0.5$, but different levels of social media biases: anti-sustainability in Fig (2(a)), neutral in Fig. (2(b)), and pro-sustainability in Fig. (2(a)).
(a) $S_{sm}=0.2$
(b) $S_{sm}=0.5$
(c) $S_{sm}=0.8$
@@ A.2.2 Social media and government feedback loops
Figure A.1: Social media feedback loop function from Eq. (3) for $\beta=0$ and different $S_{sm}$ values.
(a) $\beta=-0.15$
(b) $\beta=0.00$
(c) $\beta=+0.15$
Figure A.2: Social media feedback loop function from Eq. (3) for $S_{sm}=0.5$ and different bias values.
@@
Next, Eq. (4) ensures that an agent’s opinion changes gradually and cannot be drastically affected in a single time step. Let us delve deeper into Eq. (4). Typically, a linear equation is used to adjust an agent’s opinion based on an influence, and we plot an example of such in all upcoming figures for comparison. However, we devised a more sophisticated equation to ensure that an agent’s opinion cannot change drastically in just one step. In Eq.(4), we make use of function $f$ and $g$, which are defined in Eq. (8) and Eq. (9), respectively.
$f(A_{t-1},SM)=e^{-(1-S_{sm})\cdot|A_{t-1}-SM|}$ (8)
$g(A_{t-1},SM)=e^{S_{sm}\cdot|A_{t-1}-SM|}$ (9)
@@
Consider an agent with a susceptibility to social media of $0.5$. The two inputs that are changing are the agent’s current opinion and the social media “influence” direction, calculated one step prior using the function described in Eq. (3). Therefore, we have two inputs. To represent this in a 2D graph, we set the agent’s opinion as a parameter, set at $A_{t-1}=0.0,0.5,1.0$ and the x-axis represents the output of the social media function, indicating the opinion it promotes. The y-axis represents the agent’s updated opinion. Fig. (A.3) illustrates three functions for different values of $A_{t-1}$, and all figures contain a linear function, which is commonly used in such scenarios. Our function of choice assures that agent’s opinion cannot be changed drastically in one time step, which cannot be guaranteed using a linear function.
(a) $A_{t-1}=0.5$
(b) $A_{t-1}=0$
(c) $A_{t-1}=1$
@@
Figure A.3: Social media influence function Eq. (4) for $S_{sm}=0.5$ and different inputs of feedback
Let us begin by examining Fig. (3(a)), where the agent’s opinion set at 0.5. We can see that, no matter what opinion is promoted by social media, agent’s opinion will not change as drastically as in case of the red line. Next, let us look at Fig. (3(b)), where agent’s current opinion is at $A_{t-1}=0$. Note that with the linear function, the agent’s opinion can change from 0 to 0.4 in just one time step, which is evidently too drastic. However, with the blue function, this is not the case, ensuring a smoother and slower transition. The further away from the original opinion - the smaller the derivative value of the blue line. Additionally, this implies that agents whose opinion correspond with what they see online are more receptive to the content, accounting for a common phenomenon observed in real life - confirmation bias.
Similarly, Fig. (3(c)) illustrates the case where the agent’s opinion is set at $A_{t-1}=1$. Here, the blue line depicts the potential return, while the red line serves as a reference. In all three cases, the blue line represents the values that the agent’s opinion will take based on the input of social media influence. Notice that the derivative of the blue function decreases the further away we get from the intersection point, representing the smoother transition of opinion.
Now, let’s observe how this function changes as the susceptibility level varies. In Fig. (A.4), we plot two cases: Fig. (4(a)), where the agent’s susceptibility is $S_{sm}=0.0$, and Fig. (4(b)), where it is is $S_{sm}=1.0$. When the susceptibility level is zero, the input becomes irrelevant; social media will not influence the agent’s opinion, thus returning the same value as $A_{t-1}$, which acts as a parameter for the plot. Similarly, for a susceptibility of 1, the function becomes a linear equation, as the agent is willing to undergo drastic opinion changes in a single step.
(a) $S_{sm}=0.00$
(b) $S_{sm}=1.00$
Figure A.4: Social media influence function Eq. (4) for different levels of susceptibility.
A.3 Simulations
A.3.1 Time management
@@ A.3.3 Pseudocode
A.4 Figures
Figure A.5: Changes in average values for Model (A1) Source: our simulations.
Figure A.6: Changes in average values for Model (A2) Source: our simulations.
Figure A.7: Variances of average values for Model (A2) Source: our simulations.
Figure A.8: Changes in average values for Model (A3) Source: our simulations.
Figure A.9: Changes in average values for Model (B1) Source: our simulations.
Figure A.10: Changes in average values for Model (B2) Source: our simulations.
Figure A.11: Variances of average values for Model (B3) Source: our simulations.
Figure A.12: Changes in average values for Model (B4) Source: our simulations.
Figure A.13: Changes in average values for Model (C2) Source: our simulations.
Figure A.14: Changes in average values for Model (C3) Source: our simulations.
Figure A.15: Changes in average values for Model (C4) Source: our simulations.